OpenStudy (anonymous):

Kayla currently has an account balance of $2,134.96. She opened the account 12 years ago with a deposit of $1,327.35. If the interest compounds twice a month, what is the interest rate on the account?

OpenStudy (helder_edwin):

u have P0=1327.35 as your initial deposit. if r is the interest rate, then one term after you would have \[\large P_1=P_0+rP_0=(1+r)P_0 \] for the second term \[\large P_2=P_1+rP_1=(1+r)P_1=(1+r)(1+r)P_0=(1+r)^2P_0 \] then after k terms you would have \[\large P_k=(1+r)^kP_0 \]

OpenStudy (helder_edwin):

u have \[\large P_0=1327.35\quad\text{and}\quad P_k=2134.96 \] the question is k=? and then u solve the equation for r

OpenStudy (ranga):

Here is the compound interest formula: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\] A = Amount at maturity = $2,134.96 P = Principal Amount = $1,327.35 r = Annual interest rate in decimal = ? n = compounding period = 2 (because compounded twice a year) t = years invested = 12 Substitute and solve for r. Convert r from decimal to percentage.

OpenStudy (ranga):

Sorry n = 24 (compounded twice a month).

OpenStudy (anonymous):

okay so 2,134.96=1,327.35(1+r/2)^24

OpenStudy (ranga):

2,134.96=1,327.35(1+r/24)^288 Two changes to what you have: 1) changed r/2 to r/24 and 2) the exponent should be 288. n = 24, t = 12, nt = 24*12 = 288

OpenStudy (anonymous):

uhhh okay so it would be 2,134.96=1,327.35(1+r/24)^288? I'm not really sure how to get r though

OpenStudy (ranga):

Yeah, initially I thought it was compounded twice a year and said n = 2. But a minute later I corrected that to n = 24 in the subsequent comment.

OpenStudy (ranga):

First divide both sides by 1327.35 Then take the 288th root on both sides. That is raise both sides to ^(1/288)

OpenStudy (anonymous):

okay so 1.60843786^(1/288)=?

OpenStudy (ranga):

2,134.96 = 1,327.35 * (1+r/24)^288 2,134.96 / 1,327.35 = (1+r/24)^288 1.60843786 = (1+r/24)^288 1.60843786^(1/288) = (1 + r/24) Use your calculator or Google calculator to compute the left side.

OpenStudy (anonymous):

I got 1.00165158262

OpenStudy (ranga):

1.00165158262 = (1 + r/24) Subtract 1 from both sides. Multiply both sides by 24 to get you r

OpenStudy (ranga):

The r will be in decimal. Multiply by 100, round it off to second decimal or so and give your answer in percentage.

OpenStudy (anonymous):

i got 0.00688159425 but the only answers are a. 2.9% b. 0.3% c. 4.0% d. 4.9%

OpenStudy (ranga):

You forgot to multiply both sides by 24 to get r.

OpenStudy (anonymous):

oohhhh i divided by accident i got 3.963798288 so 4.0%?

OpenStudy (ranga):

1.60843786^(1/288) = (1 + r/24) 1.00165158262 = 1 + r/24 subtract 1 from both sides 0.00165158262 = r/24 multiply both sides by 24 0.03963798292 = r Multiply by 100 to convert to percentage r = 3.96% round it off to 4%

OpenStudy (anonymous):

okay :D thank you

OpenStudy (ranga):

you are welcome.